# Maximum-likelihood estimation of the Matérn covariance structure of isotropic spatial random fields on finite, sampled grids

### v1

@@ 4.6 Analysis of residuals
Figure 13: Maximum-likelihood analysis for the Matérn covariance structure of a thin-section of a quartzite rock under cross-polarized light [Da Mommio & Pease(2025)]. Shown are the observed data (top left) and a synthetic randomly generated from the Matérn parameters recovered as our estimate (bottom left) as discussed in Sec. 4.3, a histogram and a quantile-quantile (Q-Q) plot of the quadratic residual $2X_{\boldsymbol{\theta}}$ (middle, top), with $X_{\boldsymbol{\theta}}(\mathbf{k})$ being rendered and inspected for patterns in wave vector space (middle, bottom), as discussed in Sec. 4.6. Also shown are the expected (top right) and observed (bottom right) periodograms, $\bar{{\mathcal{S}}}_{\boldsymbol{\theta}}(\mathbf{k})$ and $|H(\mathbf{k})|^{2}$ respectively, with contour lines for the former overlain on the latter. The model fits extremely well over the entire wavenumber range.
@@
Figure 14: Maximum-likelihood analysis for the Matérn covariance structure of a patch of Venus topography. Layout and annotation are as in Fig. 13. To counter the isotropic filtering present in the original data set [Rappaport et al.(1999)Rappaport, Konopliv, Kucinskas, & Ford, Eggers(2013)], we calculated the likelihood and based our decision solely on the wavenumbers within the rendered spectral disk. The model fits very well, small residual departures visible in the Q-Q plot notwithstanding, which are indeed expected for correlated order statistics at both ends [Davison(2003)]. We draw attention to the numerical annotations in the top middle panel. The percentile corresponding to the horizontal axis limit is shown. Below that, maximum, mean, and variance of the residual $X_{\boldsymbol{\theta}}$ (see eq. 65), the test statistic $s^{2}_{X}$ (see eq. 66), the decision at the 95% significance level, and the one-sided exceedance probability for the test of it being normally distributed (see eq. 71).
@@
Figure 15: Maximum-likelihood analysis for the Matérn covariance structure of a saw-cut, polished, and sandblasted granitic rock surface fabricated in a friction experiment [Guérin-Marthe et al.(2023)Guérin-Marthe, Kwiatek, Wang, Bonnelye, Martínez-Garzón, & Dresen]. Layout and annotation are as in Figs 1314. The model fits relatively poorly under the most intransigent interpretation of the Q-Q plot, which could be argued is too demanding, as the ‘mass’ of the distribution is reasonably well aligned with statistical expectation. The residuals show relatively strong hints of unmodeled structure with directional and wavenumber dependence. On the whole, we consider this to be an acceptable model, and in future work we might be relaxing the strict adherence to eq. (71) as a make-or-break decision for model acceptance.
@@
Figure 16: Maximum-likelihood analysis for the Matérn covariance structure of Atlantic seafloor bathymetry. Layout and annotation as in Figs 1315. Of all the examples shown so far, this model fits most poorly, while it clearly still has much merit. However, its failure to be accepted based on the stringent eq. (71) will drive us to consider models beyond the isotropic stationary Matérn class. The strong directional dependence of the residuals reveals that anisotropic behavior will need to be considered, as is indeed expected both based on the intrinsic geophysical and geological properties of the seafloor [Goff & Jordan(1989b)], and from the heterogeneous and anisotropic model built [GEBCO Bathymetric Compilation Group(2019)] from compiled multiresolution data sources that include both direct shipboard seafloor sounding and interpreted satellite altimetry sea-surface measurements [Smith & Sandwell(1997)].
@@ 5 M A X I M U M - L I K E L I H O O D  P R A C T I C E
The statistical literature abounds with methods and examples of asymptotic (both in-fill and growing-domain) convergence results for specific special cases of the Matérn spectral density or covariance, that is, for specific values of $\nu$ in eqs (12) and (15), which resolve to simpler forms [Guttorp & Gneiting(2006)]. Sec. 4.3 showed how the statistical behavior of the solutions depends both on the sampling and on the random process being sampled, and it therefore must remain beyond the scope of this paper to discuss this interplay with any aspiration to generality. The important cases of anisotropy, non-stationarity, non-Gaussianity, and correlations between multivariable fields will await further treatment under our framework. Any and all of those considerations, however, will rely on a thorough understanding of the null-hypothesis of univariate, stationary, isotropic, Gaussian fields such as are provided here.
@@
For a first confrontation of our new results with geophysical and geological practice, we will limit ourselves to showing a few illuminating examples of parameter recovery, uncertainty quantification, and model testing, on completely sampled, rectangular grids. In Figs 1316, the left column contains an image of the data, underneath which we plot a simulation based on the Matérn model with the parameters derived from our procedure. The middle column appears in a layout that is a combination of the three panels of Fig. 11, showing the histogram of $X_{\boldsymbol{\theta}}(\mathbf{k})=\bar{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}$ of eq. (65) and its theoretical probability density (using the left ordinate) and quantile-quantile plot (using the right ordinate), the test statistic $s^{2}_{X}$ of eq. (66) and the decision based on it via eq. (71), below which we show the values of $X_{\boldsymbol{\theta}}(\mathbf{k})$ in wave vector space, to inspect for any undesirable patterns that may persist. The panels in the rightmost column are as the rightmost top panels in Fig. 2, showing the blurred spectral density $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$, that is, the expectation of the windowed periodogram, $\langle|H(\mathbf{k})|^{2}\rangle$, and the periodogram $|H(\mathbf{k})|^{2}$ observed from the data. Contours of the expected are drawn on the observed spectrum, highlighting where they match.
@@
The optimization was carried out by a simple unconstrained finite-difference gradient-based method (using MATLAB’s fminunc). Starting with the sample variance as a candidate for $\sigma^{2}$}, our runs where initialized with values that, while inspired by comments made by [Vanmarcke(1983)] relative to $\rho$, and [Sykulski et al.(2019)Sykulski, Olhede, Guillaumin, Lilly, & Early] relative to $\nu$, have not needed to be more than “reasonable” guesses. Up to a planar trend was removed for the data prior to analysis, and a cosine-squared window tapering 10% all around was applied to smooth any edges, exactly as we had done to make Figs 1112. See Table 1 for a summary of the results discussed next.
@@
The second example is a patch of topography from the planet Venus. Fig. 14 shows the region around Hinemoa Planitia obtained from radar altimetry mapping by the Magellan spacecraft [Rappaport et al.(1999)Rappaport, Konopliv, Kucinskas, & Ford], expanded from spherical-harmonics, reprojected to center on $(261.57^{\circ},5.29^{\circ})$, and downsampled from the original full resolution [Eggers(2013)]. The area shown occupies a mean elevation of 150 m above the mean planetary radius, with a standard deviation of 756 m. The rendered colors span the 1st through the 99th percentile, covering relative elevations 1161 m below and 2176 m above the deplaned surface. To accommodate an isotropic Nyquist filtering step applied to the original data, which effectively erased all power from the corners outside the circle inscribed in the Cartesian wave vector rectangle, we restricted the likelihood calculation to the disk-shaped domain shown in the spectral plots. The Matérn model fits the data well and is formally accepted as a null-hypothesis. Nevertheless, there are visible hints of systematic departures from complete isotropy in the middle- to large-wavenumber range that may well prove to be geologically significant upon further analysis. Furthermore, by our definition [Eggers(2013)], Hinemoa Planitia as a coherent geological unit is an irregularly bounded subdomain of the rectangular patch shown, which will require applying spatially selective windowing to fully and uniquely characterize as a locally stationary random field.
@@
The fourth and last example involves a portion of the seafloor in the North Atlantic. Fig. 16 shows the [GEBCO Bathymetric Compilation Group(2019)] bathymetric model for the area, which lies at a mean elevation of 5738 m below the sea surface, with a standard deviation of 271 m. The rendered colors span the depths between 6664 m and 4680 m. Oceanic bathymetry is a model based on few direct observations: swaths of high-resolution shipboard-derived multibeam sonar data draped over a long-wavelength low-resolution model derived from radar altimetry of the sea surface, which gravitationally reflects the topography of the ocean bottom [Smith & Sandwell(1997)]. Whether completely sampled or not, ocean bathymetry, as a mirror of plate tectonics, rarely if ever shows isotropic behavior. The isotropic Matérn model developed in this paper clearly does not apply. We include this example because our analysis allows for the testing (and rejection, as in this case) of isotropic behavior, and the initial evaluation of anisotropy, for which suitable models and estimation methods need to be developed.
Example Figure $K$ ${\sigma^{2}}/v$ $\pm$% $\nu$ $\pm$% $\pi\rho/r$ $\pm$% $\{{\sigma^{2}},\nu\}$ $\{{\sigma^{2}},\rho\}$ $\{\nu,\rho\}$
Seafloor 16 92196 0.92 3 1.26 0.6 0.01 2 38-38 +90+90 72-72
Quartzite 13 47526 0.91 4 0.91 1.0 0.02 3 31-31 +90+90 67-67
Granite 15 11556 1.64 10 1.64 1.4 0.05 4 39-39 +92+92 71-71
Venus 14 2026 1.64 53 0.31 9.0 0.34 96 22-22 +96+96 46-46
Table 1: Results from our experiments with geological and geophysical data, sorted by the degrees of freedom of the analysis, the size of the field, $K=M\times N$, adjusted to the wavenumber disk for the case of Venus. The estimate for the variance $\sigma^{2}$} is quoted as a fraction of the sample variance of the data set, here denoted $v$. In the presence of significant range, $v$ is small relative to $\sigma^{2}$}. The estimate of the range parameter is multiplied by $\pi$, see Fig. 1, and expressed as a fraction of $r$, the length of the diagonal of the data grid. For all three parameters, the one-standard-deviation estimation uncertainty is listed to the nearest per cent of the parameter itself. The final three columns contain the correlation between parameter estimates, as in Fig. 10, in per cent.
@@ 6 C O N C L U S I O N S
The [Matérn(1960)] covariance provides a parsimonious, three-parameter, characterization of stationary and isotropic geological and geophysical univariate Gaussian fields. While it has enjoyed choice application over the years, no truly pragmatic framework existed for the analysis of finite, sampled spatial data sets, including those with irregular boundaries or missing grid points. To recover the diagnostic Matérn parameters of variance, smoothness, and range, exact space-based likelihood methods are almost always impractical or computationally out of reach. Spectral-domain likelihood approximations introduced by [Whittle(1954)] hold the advantage of computational and statistical efficiency, but their parameter estimation bias has been an ongoing cause for concern, especially for incompletely and irregularly bounded sampled fields. The debiased spatial Whittle likelihood of [Guillaumin et al.(2022)Guillaumin, Sykulski, Olhede, & Simons] correctly accounts for edge effects and data omission by blurring the spectral density function in a manner that is both exact and fast. We show its efficacy in recovering Matérn covariance parameters on synthetically generated and actually observed geophysical fields. A quasi- or pseudo-likelihood, the debiased Whittle likelihood indifferently ignores wave vector correlations during parameter estimation. We show here how to calculate these interactions, exactly, and how to include them to derive the full parameter estimation covariance matrix. Only when these are properly acknowledged do we produce reliable uncertainty estimates, as we validate experimentally. Our final addition to complete the theory involves a test for the epistemic appropriateness of the Matérn model, against which possible non-stationary, an-isotropic, non-Gaussian departures from the null hypothesis can get evaluated.
Our work allows us to ask and confidently answer questions related to the statistical description of geological and geophysical fields and processes in a manner that produces solutions that are interpretable, intercomparable, and robust, with quantifiable uncertainties and tests for overall model fitness. What numbers capture the essence of a patch of spatial data? Can we estimate them, can we derive uncertainty bounds for them, can we simulate “new” realizations of fields that behave exactly as if they were derived from similar processes? The answer to all of these questions is “yes”. Our approach holds in one, two, and three dimensions, and generalizes to multiple variables, e.g. when topography and gravity are being considered jointly [[, e.g., linked by flexure, erosion, or other surface and sub-surface modifying processes, see]]Simons+2013, or for remote-sensing applications. While currently agnostic as to the application domain, our ultimate goal is to derive geophysical “process” from statistical “parameters”, e.g., to be able to assign likely formation mechanisms and histories for the data under consideration, with specific input from domain knowledge.
@@ 7 D A T A  A V A I L A B I L I T Y
All the code used to conduct the calculations and produce the figures in this paper made in Matlab by the authors is openly available as Release 2.0.0 from https://github.com/csdms-contrib/slepian_juliet, doi: 10.5281/zenodo.4085253. For an alternative code base in Python, see also [Guillaumin et al.(2025)Guillaumin, Sykulski, Olhede, & Simons]. Also compare with the R package by [Paciorek(2007)]. The quartzite photomicrograph is described by [Da Mommio & Pease(2025)] and was obtained from https://www.alexstrekeisen.it/english/meta/quartzite.php, and used with permission from Dr. Alessandro Da Mommio (alexdm83@libero.it). The original Venus topography model shtjv360 was constructed by [Rappaport et al.(1999)Rappaport, Konopliv, Kucinskas, & Ford] and resides at https://pds-geosciences.wustl.edu/missions/magellan/shadr_topo_grav/index.htm. The granite surface data pertain to the paper by [Guérin-Marthe et al.(2023)Guérin-Marthe, Kwiatek, Wang, Bonnelye, Martínez-Garzón, & Dresen] and are available from https://zenodo.org/records/6411819. The seafloor bathymetry model GEBCO_2019 is available from https://www.gebco.net/data-products/gridded-bathymetry-data/gebco-2019 as described by [GEBCO Bathymetric Compilation Group(2019)].

### v2

@@ 4.6 Analysis of residuals
Figure 13: Maximum-likelihood analysis for the Matérn covariance structure of a thin section of a quartzite rock under cross-polarized light [10]. Shown are the observed data (a) and a synthetic randomly generated from the Matérn parameters recovered as our estimate (b) as discussed in Sec. 4.3, a histogram and a quantile-quantile (Q-Q) plot of the quadratic residual $2X_{\boldsymbol{\theta}}$ (c), with $X_{\boldsymbol{\theta}}(\mathbf{k})$ being rendered and inspected for patterns in wave vector space (d), as discussed in Sec. 4.6. Also shown are the expected (e) and observed (f) periodograms, $\bar{{\mathcal{S}}}_{\boldsymbol{\theta}}(\mathbf{k})$ and $|H(\mathbf{k})|^{2}$ respectively, with contour lines (labeled by their exponent) for the former overlain on the latter. The model fits extremely well over the entire wavenumber range.
@@
Figure 14: Maximum-likelihood analysis for the Matérn covariance structure of a patch of Venus topography. Layout and annotation are as in Fig. 13. To counter the isotropic filtering present in the original data set [52, 16], we calculated the likelihood and based our decision solely on the wavenumbers within the rendered spectral disk. The model fits very well, small residual departures visible in the Q-Q plot notwithstanding, which are indeed expected for correlated order-statistics at both ends [13]. We draw attention to the numerical annotations in the top middle panel. The percentile corresponding to the horizontal axis limit is shown. Below that, maximum, mean, and variance of the residual $X_{\boldsymbol{\theta}}$ (see eq. 65), the test statistic $s^{2}_{X}$ (see eq. 66), the decision at the 95% significance level, and the one-sided exceedance probability for the test of it being normally distributed (see eq. 71).
@@
Figure 15: Maximum-likelihood analysis for the Matérn covariance structure of a saw-cut, polished, and sandblasted granitic rock surface fabricated in a friction experiment [25]. Layout and annotation are as in Figs 1314. The model fits relatively poorly under the most intransigent interpretation of the Q-Q plot, which could be argued is too demanding, as the ‘mass’ of the distribution is reasonably well aligned with statistical expectation. The residuals show relatively strong hints of unmodeled structure with directional and wavenumber dependence. On the whole, we consider this to be an acceptable model, and in future work we might relax the strict adherence to eq. (71) as a make-or-break decision for model acceptance.
@@
Figure 16: Maximum-likelihood analysis for the Matérn covariance structure of Atlantic seafloor bathymetry. Layout and annotation as in Figs 1315. Of all the examples shown so far, this model fits most poorly, while it clearly still has much merit. However, its failure to be accepted based on the stringent eq. (71) will drive us to consider models beyond the isotropic stationary Matérn class. The strong directional dependence of the residuals reveals that anisotropic behavior will need to be considered, as is indeed expected both based on the intrinsic geophysical and geological properties of the seafloor [21], and from the heterogeneous and anisotropic model [18] built from compiled multiresolution data sources that include both direct shipboard seafloor sounding and interpreted satellite altimetry sea-surface measurements [59].
@@ 5 M A X I M U M - L I K E L I H O O D  P R A C T I C E
The statistical literature abounds with methods and examples of asymptotic (infill, growing, and mixed-domain) convergence results for specific special cases of the Matérn spectral density or covariance, that is, for specific values of $\nu$ in eqs (12) and (15), which resolve to simpler forms [31]. Sec. 4.3 showed how the statistical behavior of the solutions depends both on the sampling and on the random process being sampled, and it therefore must remain beyond the scope of this paper to discuss this interplay with any aspiration to generality. The important cases of anisotropy, non-stationarity, non-Gaussianity, and correlations between multivariable fields will await further treatment under our framework. Any and all of those considerations, however, will rely on a thorough understanding of the null-hypothesis of univariate, stationary, isotropic, Gaussian fields such as are provided here.
@@
For a first confrontation of our new results with geophysical and geological practice, we will limit ourselves to showing a few illuminating examples of parameter recovery, uncertainty quantification, and model testing, on completely sampled, rectangular grids. In Figs 1316, the left column contains an image of the data, underneath which we plot a simulation based on the Matérn model with the parameters derived from our procedure. The middle column appears in a layout that is a combination of the three panels of Fig. 11, showing the histogram of $X_{\boldsymbol{\theta}}(\mathbf{k})=\bar{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(\mathbf{k})\hskip 1.00006pt|H(\mathbf{k})|^{2}$ of eq. (65), its theoretical probability density (using the left ordinate) and Q-Q plot (using the right ordinate), the test statistic $s^{2}_{X}$ of eq. (66) and the decision based on it via eq. (71). In the panel below we show the values of $X_{\boldsymbol{\theta}}(\mathbf{k})$ in wave vector space, to inspect for any undesirable patterns that may persist. The panels in the rightmost column are as the rightmost top panels in Fig. 2, showing the blurred spectral density $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$, that is, the expectation of the windowed periodogram, $\langle|H(\mathbf{k})|^{2}\rangle$, and the periodogram $|H(\mathbf{k})|^{2}$ observed from the data. Contours of the expected spectrum are drawn on the observed, highlighting where they match.
@@
The optimization was carried out by a simple unconstrained finite-difference gradient-based method (using MATLAB’s fminunc). Starting with the sample variance as a candidate for $\sigma^{2}$}, our runs were initialized with values that, while inspired by comments made by [63] relative to $\rho$, and [62] relative to $\nu$, have not needed to be more than “reasonable” guesses. Up to a planar trend was removed for the data prior to analysis, and a cosine-squared window tapering 10% all around was applied to smooth any edges, exactly as we had done to make Figs 1112. See Table 1 for a summary of the results discussed next.
@@
The second example is a patch of topography from the planet Venus. Fig. 14 shows the region around Hinemoa Planitia obtained from radar altimetry mapping by the Magellan spacecraft [52], expanded from spherical harmonics, reprojected to center on $(261.57^{\circ},5.29^{\circ})$, and downsampled from the original full resolution [16]. The area shown occupies a mean elevation of 150 m above the mean planetary radius, with a standard deviation of 756 m. The rendered colors span the 1st through the 99th percentile, which cover relative elevations 1161 m below and 2176 m above the deplaned surface. To accommodate an isotropic Nyquist filtering step applied to the original data, which effectively erased all power from the corners outside the circle inscribed in the Cartesian wave vector rectangle, we restricted the likelihood calculation to the disk-shaped domain of radius 211 km shown in the spectral plots. The Matérn model fits the data well and is formally accepted as a null-hypothesis. Nevertheless, there are visible hints of systematic departures from complete isotropy in the middle- to large-wavenumber range that may well prove to be geologically significant upon further analysis. Furthermore, by our definition [16], Hinemoa Planitia as a coherent geological unit is an irregularly bounded subdomain of the rectangular patch shown, which will require applying spatially selective windowing to fully and uniquely characterize as a locally stationary random field.
@@
The fourth and last example involves a portion of the seafloor in the North Atlantic. Fig. 16 shows the [18] bathymetric model for the area, which lies at a mean elevation of 5738 m below the sea surface, with a standard deviation of 271 m. The rendered colors span the depths between 6672 m and 4680 m. Oceanic bathymetry is a model based on few direct observations: swaths of high-resolution shipboard-derived multibeam sonar data draped over a long-wavelength low-resolution model derived from radar altimetry of the sea surface, which gravitationally reflects the topography of the ocean bottom [59]. Whether completely sampled or not, ocean bathymetry, as a mirror of plate tectonics, rarely if ever shows isotropic behavior. The isotropic Matérn model developed in this paper clearly does not apply. We include this example because our analysis allows for the testing (and rejection, as in this case) of isotropic behavior, and the initial evaluation of anisotropy, for which suitable models and estimation methods need to be developed.
Example Figure $K$ ${\sigma^{2}}/s^{2}$ $\pm$% $\nu$ $\pm$% $\rho$ $\pi\rho/r$ $\pm$% $\{{\sigma^{2}},\nu\}$ $\{{\sigma^{2}},\rho\}$ $\{\nu,\rho\}$
Seafloor 16 92196 0.92 3 1.26 0.6 1.75 km 0.01 2 38-38 +90+90 72-72
Quartzite 13 47526 0.91 4 0.91 1.0 0.0512 mm 0.02 3 31-31 +90+90 67-67
Granite 15 11556 1.64 10 1.64 1.4 0.1033 mm 0.05 4 39-39 +92+92 71-71
Venus 14 2026 1.64 53 0.31 9.0 841 km 0.34 96 22-22 +96+96 46-46
Table 1: Results from our experiments with geological and geophysical data, sorted by the degrees of freedom of the analysis, the size of the field, $K=M\times N$, adjusted to the wavenumber disk for the case of Venus. The estimate for the variance $\sigma^{2}$} is quoted as a fraction of the sample variance of the data set, here denoted $s^{2}$. In the presence of significant range, $s^{2}$ is small relative to $\sigma^{2}$}. The estimate of the range parameter $\rho$ is first presented in units of the field and then is multiplied by $\pi$, see Fig. 1, and expressed as a fraction of $r$, the length of the diagonal of the data grid. For all three parameters, the one-standard-deviation estimation uncertainty is listed to the nearest per cent of the parameter itself. The final three columns contain the correlation between parameter estimates, as in Fig. 10, in per cent.
@@ 6 C O N C L U S I O N S
The [44] covariance provides a parsimonious, three-parameter, characterization of stationary and isotropic geological and geophysical univariate Gaussian fields. While it has enjoyed choice application over the years, no truly pragmatic framework existed for the analysis of finite, sampled spatial data sets, including those with irregular boundaries or missing grid points. To recover the diagnostic Matérn parameters of variance, smoothness, and range, exact space-based likelihood methods are almost always impractical or computationally out of reach. Spectral-domain likelihood approximations introduced by [69] hold the advantage of computational and statistical efficiency, but their parameter estimation bias has been an ongoing cause for concern, especially for incompletely and irregularly bounded sampled fields. The debiased spatial Whittle likelihood of [28] correctly accounts for edge effects and data omission by blurring the spectral density function in a manner that is both exact and fast. We showed its efficacy in recovering Matérn covariance parameters on synthetically generated and actually observed geophysical fields. A quasi- or pseudo-likelihood, the debiased Whittle likelihood is indifferent to wave vector correlations during parameter estimation. We showed how to calculate these interactions, exactly, and how to include them to derive the full parameter estimation covariance matrix. Only when these are properly acknowledged do we produce reliable uncertainty estimates, as we validated experimentally. Our final addition to complete the theory involves a test for the epistemic appropriateness of the Matérn model, against which possible non-stationary, an-isotropic, non-Gaussian departures from the null-hypothesis can be evaluated.
Our work allows us to ask and confidently answer questions related to the statistical description of geological and geophysical fields and processes in a manner that produces solutions that are interpretable, intercomparable, and robust, with quantifiable uncertainties and tests for overall model fitness. What numbers capture the essence of a patch of spatial data? Can we estimate them, can we derive uncertainty bounds for them, can we simulate “new” realizations of fields that behave exactly as if they were derived from similar processes? The answer to all of these questions is “yes”. Our approach holds in one, two, and three dimensions, and generalizes to multiple variables, e.g., when topography and gravity are being considered jointly [[, e.g., linked by flexure, erosion, or other surface and sub-surface modifying processes, see]]Simons+2013, or for remote-sensing applications. While currently agnostic as to the application domain, our ultimate goal is to derive geophysical “process” from statistical “parameters”, e.g., to be able to assign likely formation mechanisms and histories for the data under consideration, with specific input from domain knowledge.
@@ 7 D A T A  A V A I L A B I L I T Y
All the code used to conduct the calculations and produce the figures in this paper made in Matlab by the authors is openly available as Release 2.0.0 from https://github.com/csdms-contrib/slepian_juliet, doi: 10.5281/zenodo.4085253. For an alternative code base in Python, see also [27]. Also see the R package by [47]. The quartzite photomicrograph is described by [10] and was obtained from https://www.alexstrekeisen.it/english/meta/quartzite.php, and used with permission from Dr. Alessandro Da Mommio (alexdm83@libero.it). The original Venus topography model shtjv360 was constructed by [52] and resides at https://pds-geosciences.wustl.edu/missions/magellan/shadr_topo_grav/index.htm. The granite surface data pertain to the paper by [25] and are available from https://zenodo.org/records/6411819. The seafloor bathymetry model GEBCO_2019 is available from https://www.gebco.net/data-products/gridded-bathymetry-data/gebco-2019 as described by [18].